OpenStudy (anonymous):

9. Theresa can ride her bike to work in 2/3 of an hour. When she drives to work it takes 1/3 of an hour. If she drives 12 miles per hour faster than she rides her bike, how fast does she ride her bike? 6 mph 12 mph 15 mph 18 mph

OpenStudy (anonymous):

we did this s = (2/3) * v s = (1/3) * (v+12) 2v = v+12 v = 12

OpenStudy (anonymous):

I remembered right after I posted.. Sorry forgot to close